OpenStudy (anonymous):

Calculate the average translational kinetic energy of a nitrogen molecule at 27 degree celcius

OpenStudy (mos1635):

CAN WE ASSUME IDEAL GAS??

OpenStudy (anonymous):

(a) Kinetic energy is a function of temperature alone: K = (3/2) · kb · T = (3/2) · 1.38065×10⁻²³J/K · 300K

OpenStudy (anonymous):

@shivam_bhalla's answer is correct, if we assume it is a ideal gas.

OpenStudy (mos1635):

yeap

OpenStudy (anonymous):

@shivam_bhalla the formula 3/2 x kb x T which formula is that and what is kb

OpenStudy (anonymous):

kb is boltzmann constant @Rizki_Zhou , yes. We assume it to be deal gas for solving this quetion

OpenStudy (anonymous):

See here for more info ---> http://hyperphysics.phy-astr.gsu.edu/hbase/kinetic/kintem.html

OpenStudy (anonymous):

thanks @shivam_bhalla

OpenStudy (anonymous):

welcome